A stone was dropped on top of the table, 2 meters above the ground. Find the distance of the ground level at time t? How long will it take the stone to reach the ground?


1
Expert's answer
2021-06-28T17:15:05-0400


Let's denote the distance from the table to the ball at time "t" as "s", and from the ball to the ground as "h". It is clear that "s + h = H" at any moment, where "H = 2m" is the height of the table.

According to the kinematic equation, have:


"s = \dfrac{gt^2}{2}"

where "g = 9.8N\/kg" is the gravitational acceleration. Thus the distance of the ground level is:


"h = H-s = H - \dfrac{gt^2}{2}"

If stone reaches the ground, "h = 0". Substituting this into the last equation, and expressing the time, obtain:


"0 = H-\dfrac{gt^2}{2}\\
\dfrac{gt^2}{2} = H\\
t = \sqrt{\dfrac{2H}{g}}\\
t = \sqrt{\dfrac{2\cdot 2}{9.8}} \approx 0.64s"

Answer. "h = H - \dfrac{gt^2}{2}", "t = 0.64s".